OpenStudy (anonymous):

place the answer in correct scientific notation, when appropriate and simplify the units. work with the units, cancel units when possible, and show simplified units in the final answer. K=1/2(3.6*10^2 kg) (2.32*10^5 m/s)^2

OpenStudy (mathstudent55):

\( K = \dfrac{1}{2} (3.6 \times 10^2 ~\text{kg})~( 2.32 \times 10^5 { ~\dfrac{\text{m}}{\text{s}} })^2 \) First, square the right quantity and remember to square the units.

OpenStudy (mathstudent55):

Are you still there?

OpenStudy (anonymous):

yea sorry i was doing something :)

OpenStudy (anonymous):

why do we square them?

OpenStudy (mathstudent55):

Isn't there an exponent of 2 for \( ( 2.32 \times 10^5 { ~\dfrac{\text{m}}{\text{s}} })^2\) ?

OpenStudy (anonymous):

53824000000 5.3*10^10 is this right?

OpenStudy (mathstudent55):

Yes, you are correct, but remember the units. Here it is, one step at a time, and with the units: First square the right quantity including the units. \( K = \dfrac{1}{2} (3.6 \times 10^2 ~kg)~( 5.38 \times 10^{10} { ~\dfrac{m^2}{s^2} }) \) Now multiply all numbers together and all units together. \( K = 9.7 \times 10^{12} ~\dfrac{kg-m^2}{s^2} \)